Exercise 17-23 (Algo) Industry Volume and Market Share Variances (LO 17-3)

D&B Ice Cream budgeted sales of 136,288 units of flavor C, assuming that the company would have 16 percent of 851,800 units sold in a particular market. The actual results were 126,630 units, based on a 14 percent share of a total market of 904,500 units. The budgeted contribution margin is $5.10 per unit.

 

Required:

Compute the sales activity variance, and break it down into market share variance and the industry volume variance. (Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option. Enter your answers rounded to the nearest whole dollar.)
Answer #1

master

16%* 851,800= 136,288

5.10* 136,288= 695,068.8


middle

16%* 904,500= 144,720

144,720* 5.10= 738,072


flex

14%* 904,500= 126,630

5.10* 126,630= 645,813


mix or market share variance: 738,072- 645,813= 92,259 U

quality or industry volume variance: 738,072- 695,068.8= 43,003.2 F

sales activity variance= 645,813- 695,068.8= 49,255.8 or 49,256 U